# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 5.2.2 Bayesian Approach
As is generally the case in supervised machine learning, there is also the variant of using Bayesian modeling for learning a target object instead of the (empirical) minimization of a loss function. To this end, one starts with a prior distribution $\rho$ over the parameter space of the reward function that is updated in light of the data set $\mathcal{D}$ by means of Bayes theorem:
$\mathbb{P}(\mathbf{\theta}\,|\,\mathcal{D})\propto L_{\mathbf{\theta}}(\mathcal{D})\cdot\rho(\mathbf{\theta})\,\text{,}$
where $L_{\mathbf{\theta}}(\mathcal{D})=\prod_{i=1}^{N}\mathbb{P}\left(\mbox{$l_{i}$ is provided}\,|\,m(q_{i}),{\mathbf{\theta}}\right)$ is the likelihood of the data under the assumed human feedback model with reward function $R_{\mathbf{\theta}}$. Such an approach is used for pairwise trajectory comparisons, for instance, by [227] for the noisy-ridge model or by [223] for the Boltzmann distribution as the human feedback model. In inverse RL, such Bayesian approaches have been considered as well (see Section 4.3 in [7]).
Instead of assuming that the reward functions are parameterized, one can use the reward functions directly as a parameter class and use a prior distribution over them. This could, for example, be a Gaussian process as initially considered by [149] for pairwise trajectory comparisons and adapted in later works [28, 60]. Here, again, it is worth mentioning that such considerations have also been made in inverse RL before (see Section 4.3 in [7]).
5.2.3 Direct Methods
Note that the two-step approach of utility learning and policy optimization is not the only viable path to learning a policy from human feedback. While the previous two subsections have discussed the case in which we learn a reward function from observed preferences by assuming a human feedback model, an emerging branch of the literature is concerned with circumventing the reward-learning step and using preferences directly to address the actual RL problem. Concrete approaches are DPO [213], SLiC-HF [292], OPPO [134], or DPPO [6]. [11] introduce an objective called $\Psi$-preference optimization ($\Psi$PO) that unifies the objective functions in DPO and RLHF. More specifically, for a specific instantiation of $\Psi$, the objective in $\Psi$PO recovers the latter two.
It is worth noting that approaches for directly learning the policy from preferences have been considered in the past as well [269, 83, 271, 272, 41]. In Sections 3.2.1 and 3.2.2 in the survey by [270], these methods are explained in more detail.
@@ 5.2.4 Partial Identifiability
A crucial question when it comes to learning the reward function is whether the reward function can be identified at all. If two reward functions induce exactly the same human feedback model, the reward function is called partially identifiable or ambiguous. [241] study this topic for the Boltzmann distribution as the underlying human feedback model when demonstrations (inverse RL) or pairwise trajectory preferences are given as feedback. For demonstrations, this question has also been examined in other works [195, 71, 139, 43].
@@ 5.2.5 Interpretability
The field of explainable artificial intelligence (XAI) has emerged in recent years to improve the transparency and explainability of models or even to enable them in the first place. Roughly speaking, the aim is to resort to more interpretable methods or provide explanations for both experts and non-experts, shedding light on why a certain input in a (black box) model leads to a certain result. Explanations can take different forms, as can the ways to ensure the transparency of models, and for a detailed overview, we refer to [17]. It is worth noting that the field has grown so extensively over the years that even dedicated overviews for the field of interpretable and explainable RL are by now available [209, 97, 211, 183].
For the branch of RLHF, the existing works are quite sparse and mostly limited to using tree models as transparent and explainable models for learning the reward function [26, 25, 132, 24, 133]. Another way to realize explainability within RLHF suggested by [289] is to learn simultaneously the reward function and the importance of states using a weight network. Assuming that for (long) trajectories, only a few states are important for the preference outcome, their framework can be used to select samples for explainability purposes. Moreover, a perturbation analysis is suggested to evaluate explanations in a quantitative manner using the learned state importance weights.
@@ 5.2.6 Online Improvements
[58] demonstrate that it is important to improve the reward model online, which is also confirmed by [86] as otherwise, the issue of overoptimization of an imperfect reward model may occur. [2] give an example of this: They attempt to fine-tune an agent initialized with behavioral cloning with an engineered reward function and find that it fails to generalize and actually worsens the performance. They also compare RLHF with a reward model trained offline with iterative improvement and find that this leads to better performance, even sometimes exceeding human performance. However, this is not generally the case, as [178] show that reward models trained online together with a policy may not be effective when a completely new policy is trained.
5.2.7 Combining Different Feedback Types
An additional extension that is of interest is how the learning process can incorporate different sources of feedback. So far, three options have been considered in the literature: First, the combination of demonstrations and preferences in a model-free approach [119] or a passive-active preference-based reward learning approach [202, 29]. Second, the combination of pairwise preferences and ordinal labels for actions in order to ensure safety-aware learning [60]. And third, the combination of pairwise preferences, demonstrations, and ratings for reducing the cognitive load for the human labeler  [146].
@@ 5.3 Evaluating Learned Reward Functions
A central question when it comes to learning the reward function is how to evaluate the learned reward function and how reward functions can be compared with each other. For this purpose, a couple of different approaches are available.
@@
In inverse RL, a common method for evaluation is the rollout method [194, 80, 119, 40]. In this approach, one first learns an optimal policy for the learned reward function and then estimates the value of this policy for online trajectory rollouts using the known ground-truth reward function. This approach can be transferred to RLHF as well. In many cases, however, the true reward function is unknown, and in safety-critical areas such as medicine or autonomous driving, such online rollouts cannot be executed.
@@
As a possible alternative, so-called off-policy evaluations, which estimate the value of the optimal policy on the basis of an available data set, have been established. For coping with biases or large variances due to policy mismatch, approaches using importance sampling [208], regression- or classification-based methods [201, 153, 120], or combinations of these [128] have been proposed. The problem with these approaches, however, is that the traces of the explicit sources of error through policy learning or reward learning are blurred.
@@
Yet another alternative, which has been advanced in the seminal paper by [98], is using a suitable distance function for reward functions. Suitable here means that two reward functions, which differ only by certain transformations such as potential shaping [195] or positive scaling, should have zero distance if these transformations do not change the policy ranking with regard to the expected return. For this purpose, [98] present a pseudometric, called Equivalent-Policy Invariant Comparison (EPIC) distance, that is determined in three steps: First, mapping two reward functions to a so-called canonicalization form that is invariant to transformations of the latter kind. Second, normalizing these canonicalization forms by means of a specific weighted $L_{2}$-norm whose weights are determined by a distribution over the transitions. Finally, the EPIC distance is the weighted $L_{2}$-norm distance of the normalized canonicalization forms.
Even if some attractive properties, above all a Lipschitz continuity in terms of the EPIC distance of two reward functions for the difference of the value functions of the induced optimal policies is shown, this distance has its shortcomings. One of these is that the canonicalization form used by EPIC distance does not encode sufficient knowledge about the underlying transition dynamics, which might lead to unreliable distances when evaluating reward functions on physically non-realizable transitions. To this end, [276] propose the Dynamics-Aware Reward Distance (DARD), which uses a slightly different form of canonicalization but restricts the evaluation of the reward functions to transitions that are approximately physically feasible.
Recently, EPIC-like distances [124] and STAndardised Reward Comparison (STARC) metrics [240], which are entire classes of pseudometrics on the space of all reward functions were proposed that generalize the three-step approach underlying the EPIC distance (and DARD) by parameterizing each of the steps. Specifically, the canonicalization function in the first step, the normalization in the second, and the metric in the third step are kept variable. If these three functional parameters fulfill certain requirements, then the resulting distance has some appealing properties, e.g., being a pseudometric that is zero if and only if the two reward functions induce the same ordering of policies or imply upper and lower bounds on value function differences.
@@
For an evaluation by visual inspection, [123] propose a method for preprocessing reward functions by transforming them into simpler but equivalent reward functions for better interpretability. Related to this and the rollout method, the quality of the reward function learned can also be evaluated by a human (or expert) by examining the behavior of the agent on the target task.
@@ 5.4 Reward Model Inputs
For instance, many typical RL benchmarks are in the continuous control domain (e.g., MuJoCo simulated robotics tasks) with true state representations and simple discrete actions. In such problems, [58] train reward models from these inputs.
When no compact state representation is available, raw images are often used in control tasks, which makes the learning of rewards more challenging since the setting becomes partially observable and the reward function is generally not Markov with respect to the observations. In such cases, the conventional trick of approximating a true state with a sequence of frames is often employed. This approach is used, for instance, by [58] to train reward models on the Atari benchmark suite.
More recently, many applications of RLHF are in the natural language processing (NLP) domain. In these settings, the policy takes natural language as both input and output while the reward model takes it as input (e.g., [200]). Naturally, more complex scenarios (e.g., with both language and vision inputs [2]) have also been studied.
@@ 5.5.1 Using Prior Data
There are often large amounts of prior data available at little or no additional cost. While this data generally was generated for other tasks, many basic human preferences are the same for various tasks and can often even be extracted from completely unsupervised data such as text corpora. By leveraging this prior data, we can greatly reduce the amount of feedback necessary to learn the current task’s objective. We explore various methods, including foundation models, meta- and transfer learning, pertaining, and reward model initialization.
@@
Meta- and transfer learning techniques in reward model training exploit the commonalities across different objectives, facilitating quick adaptation to new tasks. [219] develop a broadly applicable meta-reward model, pre-trained on a diverse set of tasks to capture a wide range of preference patterns, enabling efficient adaptation to new tasks with fewer examples. [278] use a similar meta-learning approach to build a goal classifier across multiple visuomotor tasks. Closely related to these meta-learning approaches, [113] integrate few-shot learning principles, optimizing their approach for scenarios where only a few examples are available for adapting to new tasks. In the domain of transfer learning, [169] explore zero-shot transfer of preferences, a method that enables adapting preferences without additional data from the new task. In a different vein, but closely related to meta- and transfer learning, [179] tackle the lifelong inverse RL problem, focusing on inferring reward functions for multiple tasks over time, which involves knowledge transfer between tasks. Collectively, these studies underscore the potential of meta- and transfer learning in enhancing the efficiency and applicability of reward models in RLHF.
@@
Foundation models, i.e., models trained on large amounts of often unlabeled data, can acquire significant knowledge about basic human preferences. A language model trained to predict the next token in a text corpus, for example, may learn to complete the sentence ‘Frank was mad that his vacuum robot broke the vase,’ thereby learning that humans prefer non-destructive behavior. These learned preferences can then be leveraged in RLHF approaches. [70] exemplify such an approach by training a success detector using a pre-trained vision-language model (Flamingo). Their approach utilizes a data set of trajectories with binary success labels, employing a non-interactive training method.
@@
It is often beneficial to initialize the reward model with parameters from a model trained on a related task. This strategy is particularly common in language model fine-tuning, where self-supervised pretraining is a common practice. In such scenarios, it becomes logical to use these pre-trained models for initializing not just the policy but also the reward model. This methodology is adopted by [9] and [200]. Specifically, [200] use a pretrained language model for the reward model, opting for a smaller model relative to the policy to mitigate unstable learning. Notably, while they apply supervised fine-tuning to the policy before the RLHF phase, the reward model is initialized directly from the language model without any preliminary fine-tuning. This approach’s applicability extends beyond language models to other areas. A notable example is [2], who, in the control domain, begin by training a policy through contrastive self-supervised learning and behavioral cloning. They then add an MLP head to the policy for the prediction of cumulative rewards.
Preference Model Pretraining
Preference-model pretraining [9, 12] leverages prior offline data to pretrain the preference model before training it on policy samples. [9] note that in the case of language models, noisy preference data can be readily obtained from sources such as rated Reddit comments, preferred Stack Overflow answers, and reverted Wikipedia edits. They leverage this as a pretraining step to increase data efficiency. This is in addition to regular language model pretraining, as discussed in the previous paragraph. A similar approach could be applied to control in case prior data and some means of inferring preferences, such as human corrections, are available. Even if no inferred preferences are available, [259] show that it can be beneficial to pre-train the preference model to predict close to constant reward on an initial set of trajectories. This avoids excessive fitting of the policy to random initialization differences in the reward function.
@@
A compact representation that captures all relevant information for human preferences while minimizing noise can greatly enhance preference learning efficiency. It may also generalize better than a representation learned end-to-end as part of the preference learning task, which may contain spurious correlations. [36] address this by proposing the learning of features through explicit human feedback using feature traces. Feature traces (see Section 3.2.9) involve human labelers explicitly teaching relevant features one by one by demonstrating behavior in which the feature monotonically increases or decreases. This method directly aligns the learned representation with human-identified features, enhancing preference learning efficiency but requiring detailed human input. However, feature traces require labelers to be able to identify and articulate relevant features, which can be challenging. [34] offer an alternative approach with their Similarity-based Implicit Representation Learning (SIRL) method. SIRL learns representations from similarity queries (see Section 3.2.10), where human labelers provide feedback on whether behaviors are similar or different concerning the features that matter to them. This method captures a broader range of human notions of similarity without needing explicit feature knowledge, thus reducing the cognitive load on human labelers. In summary, while both approaches emphasize human feedback’s centrality in representation learning, they differ in their methods of gathering this feedback. The feature traces used by [36] require specific feature knowledge, whereas SIRL used by [34] utilizes more intuitive similarity assessments, potentially offering a more user-friendly way to capture human preferences.

### v2

@@ 5.2.2 Bayesian Approach
As is generally the case in supervised ML, there is also the variant of using Bayesian modeling for learning a target object instead of the (empirical) minimization of a loss function. To this end, one starts with a prior distribution $\rho$ over the parameter space of the reward function that is updated in light of the data set $\mathcal{D}$ by means of Bayes theorem:
$\mathbb{P}(\bm{\psi}\,|\,\mathcal{D})\propto L_{\bm{\psi}}(\mathcal{D})\cdot\rho(\bm{\psi})\,\text{,}$
where $L_{\bm{\psi}}(\mathcal{D})=\prod_{i=1}^{N}\mathbb{P}\left(\mbox{$l_{i}$ is provided}\,|\,m(q_{i}),{\bm{\psi}}\right)$ is the likelihood of the data under the assumed human feedback model with reward function $R_{\bm{\psi}}$. Such an approach is used for pairwise trajectory comparisons, for instance, by Schoenauer et al. (2014) for the noisy-ridge model or by Sadigh et al. (2017) for the Boltzmann distribution as the human feedback model. In inverse RL, such Bayesian approaches have been considered as well (see Section 4.3 in Arora & Doshi (2021)).
Instead of assuming that the reward functions are parameterized, one can use the reward functions directly as a parameter class and use a prior distribution over them. This could, for example, be a Gaussian process as initially considered by Kupcsik et al. (2018) for pairwise trajectory comparisons and adapted in later works (Bıyık et al., 2020; Cosner et al., 2022). Here, again, it is worth mentioning that such considerations have also been made in inverse RL before (see Section 4.3 in Arora & Doshi (2021)).
@@ 5.2.3 Partial Identifiability
A crucial question when it comes to learning the reward function is whether the reward function can be identified at all. If two reward functions induce exactly the same human feedback model, the reward function is called partially identifiable or ambiguous. Skalse et al. (2023) study this topic for the Boltzmann distribution as the underlying human feedback model when demonstrations (inverse RL) or pairwise trajectory preferences are given as feedback. For demonstrations, this question has also been examined in other works (Ng & Russell, 2000; Dvijotham & Todorov, 2010; Kim et al., 2021; Cao et al., 2021a). On a related note, Ellis et al. (2024) tackle this identifiability issue by considering suitable acquisition functions (see Section 4.1.1).
@@ 5.2.4 Interpretability
The field of explainable artificial intelligence (XAI) has emerged in recent years to improve the transparency and explainability of models or even to enable them in the first place. Roughly speaking, the aim is to resort to more interpretable methods or provide explanations for both experts and non-experts, shedding light on why a certain input in a (black box) model leads to a certain result. Explanations can take different forms, as can the ways to ensure the transparency of models, and for a detailed overview, we refer to Barredo Arrieta et al. (2020). It is worth noting that the field has grown so extensively over the years that even dedicated overviews for the field of interpretable and explainable RL are by now available (Puiutta & Veith, 2020; Glanois et al., 2022; Qing et al., 2023; Milani et al., 2023).
For the branch of RLHF, the existing works are quite sparse and mostly limited to using tree models as transparent and explainable models for learning the reward function (Bewley & Lécué, 2022; Bewley et al., 2022; Kalra & Brown, 2022; Bewley et al., 2024; Kalra & Brown, 2023). Another way to realize explainability within RLHF suggested by Zhang & Kashima (2023) is to learn simultaneously the reward function and the importance of states using a weight network. Assuming that for (long) trajectories, only a few states are important for the preference outcome, their framework can be used to select samples for explainability purposes. Moreover, a perturbation analysis is suggested to evaluate explanations in a quantitative manner using the learned state importance weights.
@@ 5.2.5 Online Improvements
Christiano et al. (2017) demonstrate that it is important to improve the reward model online, a finding that has been confirmed by subsequent works such as the one by Gao et al. (2023), which empirically demonstrates that overoptimization of a reward model trained offline leads to performance degradation. Without online improvements, issues of overoptimization of an imperfect reward model may occur. Abramson et al. (2022) give an example of this: They attempt to fine-tune an agent initialized with behavioral cloning with an engineered reward function and find that it fails to generalize and actually worsens the performance. They also compare RLHF with a reward model trained offline with iterative improvement and find that iterative improvement leads to better performance, even sometimes exceeding human performance.
This is related to issues posed by the approximate nature of the reward model in general, discussed in further detail in Section 6.1, but improving reward model accuracy, in general, is not sufficient: McKinney et al. (2022) further show the interdependence of the reward model and the policy, demonstrating that reward models trained online together with a policy may not be effective when a completely new policy is trained.
Solutions to the problems of overoptimization and interdependence can take different forms: One is to update the reward model online with sufficient frequency using notably more on-policy queries (see Section 4.1.1), another is to improve the reward model, e.g., by leveraging ensembles (Coste et al., 2024) or by modifying the training procedure to place additional emphasis on challenging examples (Zheng et al., 2024), and a third, discussed in Section 6.1, is to add constraints to the policy training.
5.2.6 Learning from Multiple Feedback Types
As discussed in Section 3.5, it is often desirable to combine several feedback types. This requires extensions of the learning process to incorporate different sources of feedback. Learning from multiple feedback types can be achieved by pre-processing the feedback, assuming common latent factors, or by using the feedback types for distinct purposes.
The first approach is demonstrated by (Novoseller et al., 2023), who infer preferences from demonstrations, allowing them to treat both types of feedback equally in the learning pipeline. In the style of the second approach, Jeon et al. (2020) proposes the unified framework of reward-rational choice, which allows for interpreting many forms of human feedback as Boltzmann-rational choices and, through this common framework, enables combination and adaptive selection of feedback types. Finally, different types of feedback can be used for entirely different purposes, such as one for objective learning and another for safety-constraints (Cosner et al., 2022) or for representation learning (see Section 3.2.3).
Since multiple sources of reward information may conflict, it is important to consider how to combine them. Krasheninnikov et al. (2021) study several possible strategies of combining several reward functions in this setting, relating it to multi-task inverse RL. Note that this challenge of conflicting sources of reward relates tightly to challenges posed when receiving diverse preferences from different labelers, as discussed in Section 5.1.6.
5.2.7 Offline Reward Learning
There is a recent trend towards offline RLHF, where both the reward model and the policy are trained offline. The offline setting is also frequently considered in RLHF theory (Section 7). Early approaches in this area (Kim et al., 2023; Shin et al., 2023) first generate queries from an offline dataset of behaviors, gather human responses, train a reward model from the resulting preference data, and then leverage offline RL algorithms to derive a policy. We do not cover these works in detail, since this survey primarily focuses on the interactive and online setting (see Section 1.3). Nonetheless, the offline setting is particularly useful for evaluating novel approaches, e.g., for active query selection, using offline datasets. We refer the interested readers to Section 8.4 for a discussion of available datasets.
@@ 5.3 Evaluating Learned Reward Functions
A central question when it comes to learning the reward function is how to evaluate the learned reward function and how reward functions can be compared with each other. For this purpose, different approaches are available, e.g.:
@@
In inverse RL, a common method for evaluation is the rollout method (Ng et al., 1999; Fu et al., 2018a; Ibarz et al., 2018; Brown et al., 2019). In this approach, one first learns an optimal policy for the learned reward function and then estimates the value of this policy for online trajectory rollouts using the known ground-truth reward function. This approach can be transferred to RLHF as well. In many cases, however, especially in safety-critical areas such as medicine or autonomous driving, such online rollouts cannot be executed.
@@
When online rollouts are not possible, so-called off-policy evaluations, which estimate the value of the optimal policy on the basis of an available data set, may be considered. For coping with biases or large variances due to policy mismatch, approaches using importance sampling (Precup et al., 2000), regression- or classification-based methods (Paduraru, 2013; Le et al., 2019; Irpan et al., 2019), or combinations of these (Jiang & Li, 2016) have been proposed. The problem with these approaches, however, is that the traces of the explicit sources of error through policy learning or reward learning are blurred, and that they require access to the ground-truth rewards.
@@
Yet another alternative, which has been advanced in the seminal paper by Gleave et al. (2022a), is using a suitable distance function for reward functions. Suitable here means that two reward functions, which differ only by certain transformations such as potential shaping (Ng & Russell, 2000) or positive scaling, should have zero distance if these transformations do not change the policy ranking with regard to the expected return. For this purpose, Gleave et al. (2022a) present a pseudometric, called Equivalent-Policy Invariant Comparison (EPIC) distance, that is determined in three steps: First, mapping two reward functions to a so-called canonicalization form that is invariant to transformations of the latter kind. Second, normalizing these canonicalization forms by means of a specific weighted $L_{2}$-norm whose weights are determined by a distribution over the transitions. Finally, the EPIC distance is the weighted $L_{2}$-norm distance of the normalized canonicalization forms.
Even if some attractive properties, above all a Lipschitz continuity in terms of the EPIC distance of two reward functions for the difference of the value functions of the induced optimal policies is shown, this distance has its shortcomings. One of these is that the canonicalization form used by EPIC distance does not encode sufficient knowledge about the underlying transition dynamics, which might lead to unreliable distances when evaluating reward functions on physically non-realizable transitions. To this end, Wulfe et al. (2022) propose the Dynamics-Aware Reward Distance (DARD), which uses a slightly different form of canonicalization but restricts the evaluation of the reward functions to transitions that are approximately physically feasible.
Recently, EPIC-like distances (Jenner et al., 2022) and STAndardised Reward Comparison (STARC) metrics (Skalse et al., 2024), which are entire classes of pseudometrics on the space of all reward functions were proposed that generalize the three-step approach underlying the EPIC distance (and DARD) by parameterizing each of the steps. Specifically, the canonicalization function in the first step, the normalization in the second, and the metric in the third step are kept variable. If these three functional parameters fulfill certain requirements, then the resulting distance has some appealing properties, e.g., being a pseudometric that is zero if and only if the two reward functions induce the same ordering of policies or imply upper and lower bounds on value function differences. In particular, these metrics retain the flexibility of DARD (in terms of specifying transition dynamics), while at the same time preserving the theoretical justification of EPIC.
@@
For an evaluation by visual inspection, Jenner & Gleave (2021) propose a method for preprocessing reward functions by transforming them into simpler but equivalent reward functions for better interpretability. Related to this and the rollout method, the quality of the reward function learned can also be evaluated by a human (or expert) by examining the behavior of the agent on the target task. Besides, in the context of LLMs, datasets have been proposed and specifically designed to evaluate the (in)consistency of learned reward models with respect to semantic changes of prompts (Shen et al., 2024).
@@ 5.4 Reward Model Inputs
For instance, many typical RL benchmarks are in the continuous control domain (e.g., MuJoCo simulated robotics tasks) with true state representations and simple discrete actions. In such problems, Christiano et al. (2017) train reward models from these inputs.
When no compact state representation is available, raw images are often used in control tasks, which makes the learning of rewards more challenging since the setting becomes partially observable and the reward function is generally not Markov with respect to the observations. In such cases, the conventional trick of approximating a true state with a sequence of frames is often employed. This approach is used, for instance, by Christiano et al. (2017) to train reward models on the Atari benchmark suite. When taking only observations as inputs, one can resort to recurrent models (Abramson et al., 2022) or Transformer-based models (Kim et al., 2023).
More recently, many applications of RLHF are in the natural language processing (NLP) domain. In these settings, the policy takes natural language as both input and output while the reward model takes it as input (see, e.g., the work by Ouyang et al. (2022)). Naturally, more complex scenarios (e.g., with both language and vision inputs (Abramson et al., 2022)) have also been studied.
@@ 5.5.1 Using Prior Data
There are often large amounts of prior data available at little or no additional cost. While this data generally was generated for other tasks, many basic human preferences are the same for various tasks and can often even be extracted from completely unsupervised data such as text corpora. By leveraging this prior data, we can greatly reduce the amount of feedback necessary to learn the current task’s objective. We explore various methods, including meta- and transfer learning, leveraging foundation models, reward model initialization, preference model pretaining, and supervised representation learning.
@@
Meta- and transfer learning techniques in reward model training exploit the commonalities across different objectives, facilitating quick adaptation to new tasks. Ren et al. (2022) develop a broadly applicable meta-reward model, pre-trained on a diverse set of tasks to capture a wide range of preference patterns, enabling efficient adaptation to new tasks with fewer examples. Xie et al. (2018) use a similar meta-learning approach to build a goal classifier across multiple visuomotor tasks. Closely related to these meta-learning approaches, Hejna & Sadigh (2022) integrate few-shot learning principles, optimizing their approach for scenarios where only a few examples are available for adapting to new tasks. In the domain of transfer learning, Liu et al. (2023a) explore zero-shot transfer of preferences, a method that enables adapting preferences without additional data from the new task. In a different vein, but closely related to meta- and transfer learning, Mendez et al. (2018) tackle the lifelong inverse RL problem, focusing on inferring reward functions for multiple tasks over time, which involves knowledge transfer between tasks. Collectively, these studies underscore the potential of meta- and transfer learning in enhancing the efficiency and applicability of reward models in RLHF.
@@
Foundation models, i.e., large models trained on large amounts of often unlabeled data, can acquire significant knowledge about basic human preferences. A language model trained to predict the next token in a text corpus, for example, may learn to complete the sentence ‘Frank was mad that his vacuum robot broke the vase,’ thereby learning that humans prefer non-destructive behavior. These learned preferences can then be leveraged in RLHF approaches. For instance, Kwon et al. (2023) propose to use LLM as a source of rewards. Du et al. (2023) is another example, where a success detector is trained using a pre-trained vision-language model (Flamingo). Their approach utilizes a data set of trajectories with binary success labels, employing a non-interactive training method.
@@
It is often beneficial to initialize the reward model with parameters from a model trained on a related task. This strategy is particularly common in language model fine-tuning, where self-supervised pretraining is a common practice. In such scenarios, it becomes logical to use these pre-trained models for initializing not just the policy but also the reward model. This methodology is adopted by Askell et al. (2021) and Ouyang et al. (2022). Specifically, Ouyang et al. (2022) use a pretrained language model for the reward model, opting for a smaller model relative to the policy to mitigate unstable learning. Notably, while they apply supervised fine-tuning to the policy before the RLHF phase, the reward model is initialized directly from the language model without any preliminary fine-tuning. This approach’s applicability extends beyond language models to other areas. A notable example is Abramson et al. (2022), who, in the control domain, begin by training a policy through contrastive self-supervised learning and behavioral cloning. They then add an MLP head to the policy for the prediction of cumulative rewards.
Reward Model Pretraining
Reward model pretraining (Askell et al., 2021; Bai et al., 2022a) leverages prior offline data to pretrain the preference model before training it on policy samples. Askell et al. (2021) note that in the case of language models, noisy preference data can be readily obtained from sources such as rated Reddit comments, preferred Stack Overflow answers, and reverted Wikipedia edits. They leverage this as a pretraining step to increase data efficiency. This is in addition to regular language model pretraining, as discussed in the previous paragraph. A similar approach could be applied to control in case prior data and some means of inferring preferences, such as human corrections, are available. Even if no inferred preferences are available, Verma & Kambhampati (2023a) show that it can be beneficial to pre-train the preference model to predict close to constant reward on an initial set of trajectories. This avoids excessive fitting of the policy to random initialization differences in the reward function.
@@
A compact representation that captures all relevant information for human preferences while minimizing noise can greatly enhance preference learning efficiency. It may also generalize better than a representation learned end-to-end as part of the preference learning task, which may contain spurious correlations. Bobu et al. (2022) address this by proposing the learning of features through explicit human feedback using feature traces. Feature traces (see Section 3.2.3) involve human labelers explicitly teaching relevant features one by one by demonstrating behavior in which the feature monotonically increases or decreases. This method directly aligns the learned representation with human-identified features, enhancing preference learning efficiency but requiring detailed human input. However, feature traces require labelers to be able to identify and articulate relevant features, which can be challenging. Bobu et al. (2023) offer an alternative approach with their Similarity-based Implicit Representation Learning (SIRL) method. SIRL learns representations from similarity queries (see Section 3.2.3), where human labelers provide feedback on whether behaviors are similar or different concerning the features that matter to them. This method captures a broader range of human notions of similarity without needing explicit feature knowledge, thus reducing the cognitive load on human labelers. In summary, while both approaches emphasize human feedback’s centrality in representation learning, they differ in their methods of gathering this feedback. The feature traces used by Bobu et al. (2022) require specific feature knowledge, whereas SIRL used by Bobu et al. (2023) utilizes more intuitive similarity assessments, potentially offering a more user-friendly way to capture human preferences.